OpenStudy (anonymous):

please help they ask me What is the slope of the line which passes through (–6, 0) and (4, 0)? and it gave me undefine can someone check if am right?

OpenStudy (lgbasallote):

check if you're right? your answer is undefined?

OpenStudy (anonymous):

undefined –2 0 2 those are the choices?

OpenStudy (lgbasallote):

which one is *your* answer?

OpenStudy (anonymous):

the first on XD

OpenStudy (lgbasallote):

let's check \[\text{slope} = \frac{y_2-y_1}{x_2-x_1}\] \[\text{slope} = \frac{0 - 0}{4 - (-6)}\] is this what you did?

OpenStudy (anonymous):

YEAP

OpenStudy (anonymous):

(-1,0)

OpenStudy (lgbasallote):

what's (-1,0)?

OpenStudy (anonymous):

way I did something wrong

OpenStudy (anonymous):

0/10

OpenStudy (lgbasallote):

yup so what's that equal to?

OpenStudy (anonymous):

undifined :D

OpenStudy (anonymous):

I AM NOT SURE :(

OpenStudy (anonymous):

well if its 0 then its undifined

OpenStudy (lgbasallote):

\[\huge \frac A0 = \text{undefined}\] \[\huge \frac 0A = 0\]

OpenStudy (anonymous):

oh!!!! I thought that when ever you had zero it was undefined

OpenStudy (anonymous):

ohhh!!!

OpenStudy (lgbasallote):

haha ironic isnt it...back in middle school people memorize that 0 divided by anything is 0...then undefined is introduced and they get confused :)

OpenStudy (anonymous):

hehe right! thank you so much!

OpenStudy (lgbasallote):

<tips imaginary hat>

OpenStudy (anonymous):

haha